# Eternal Universes

### v1

@@ Abstract
We consider the possibility of a past and future eternal universe, constructing geodesically complete inflating, loitering, and bouncing spacetimes. We identify the constraints energy conditions in General Relativity place on the building of eternal cosmological models. Inflationary and bouncing behavior are shown to be essential ingredients in all significant examples. Non-trivial complete spacetimes are shown to violate the null energy condition (NEC) for at least some amount of time, although, some obey the average null energy condition. Ignoring the intractable subtleties introduced by quantum considerations, such as rare tunneling events and Boltzmann brains, we demonstrate that these universes need not have had a beginning or an end.
@@ I Introduction
Whether or not the universe had a beginning is a question that naturally arises within our most successful theory of the early universe, the theory of inflation Guth (1981); Linde (1982); Albrecht and Steinhardt (1982). Soon after the theory’s inception it became clear that once inflation started, it would never stop. While early-time inflation has ceased in our observable universe, allowing galaxies and structures to form, it continues indefinitely in other parts of the universe, and hence, inflation is eternal into the future. Specifically, quantum fluctuations in the field responsible for the accelerated expansion can cause the field to remain in its high-energy state, sustaining inflation in those regions indefinitely 198 (1983); Vilenkin (1983); Linde (1986); Guth (2007). In a Letter by Borde, Guth and Vilenkin (BGV) Borde et al. (2003), it was subsequently argued that inflation cannot likewise be eternal into the past; although, we have countered this belief Easson and Lesnefsky (2024); Lesnefsky et al. (2023a) and present further evidence that inflation can, in fact, be eternal into the past in the current work. Acknowledging the possibility of a past eternal universe opens entirely new avenues of research. We show there are multiple ways that the universe can sustain itself within classical General Relativity for an eternity, albeit at the expense of violation of certain energy conditions.
Three broad categories of an eternal universe are suggested. Our examples are admittedly toy models; however, given the wide variety we expect our findings to be quite general and applicable to many solutions of interest. In particular, we will construct eternal inflationary, loitering, and bouncing cosmologies. In each case we will provide a rigorous proof of full geodesic completeness. Existing proposed no-go theorems are circumvented or obviated as discussed in detail in an accompanying Letter Easson and Lesnefsky (2024). We show that all (non-trivial) models that are eternal geodesically complete necessarily involve a phase of accelerated expansion (or inflationary phase) for at least some amount of time and generically include a bouncing phase in accordance with our reported definitions.
The models we present violate the traditional energy conditions, a property we will conjecture is necessary in GR for geodesic completeness.  [^3] A contained discussion of energy conditions is offered in Appendix A. In some instances violations of the null energy condition (NEC) can be limited to an arbitrarily short amount of time $\Delta t$, and confined to the primordial past.  [^4] In its most provocative form, the Heisenberg uncertainty principle is written $\Delta E\Delta t\geq\hbar/2$. If the time interval is made very small the uncertainty principle may result in a very large energy fluctuation which can back-react on the spacetime. Models presented here having such behavior are capable of satisfying the average null energy condition (ANEC).
@@
[…]
where the integral is over a null geodesic $\gamma$ with tangent vector $k^{\mu}=dx^{\mu}/d\lambda$, and $\lambda$ is an affine parameter with respect to which the tangent vector to the geodesic is defined. In GR, it is frequently convenient to replace $T_{\mu\nu}$ in Eq. 1.1 by the Ricci tensor $R_{\mu\nu}$, simply using the Einstein field equations and the definition of a null vector. The ANEC has been shown to hold even in the case of arbitrary Casimir systems as long as the geodesic does not intersect or asymptotically approach the plates Graham and Olum (2005); Fewster et al. (2007). If the null geodesic is not assumed to be achronal, the ANEC may be violated by a quantum scalar field in a spacetime compactified in one spatial dimension or in the spacetime around a Schwarzschild black hole Visser (1996). As we shall see some of the eternal models we construct are capable of satisfying the ANEC.
@@
This is the companion paper to our Letter Easson and Lesnefsky (2024). A few of the Theorems and Corollaries we discuss were presented without proof in our earlier work Lesnefsky et al. (2023b); Easson and Lesnefsky (2024). The proofs are revealed presently.
An outline of this paper is as follows. In section II, we prove a theorem which definitively provides the criteria for geodesic completeness of generalized Friedmann-Robertson-Walker (GFRW) spacetimes. In section III, we discuss the definition of inflation and bouncing cosmologies and provide an example of an eternally inflating universe which is eternal and geodesically complete both into the past and future. In section IV, we provide examples of eternal bouncing cosmologies which are geodesically complete both into the past and future. In section V, we provide examples of loitering (and quasi-loitering) models which are eternal and geodesically complete both into the past and future. Finally, in section VI, we present a discussion and our conclusions. We provide appendices to discuss the energy conditions in General Relativity with respect to cosmology, to provide proof details and to define Generalized Friedmann-Robertson-Walker (GFRW) spacetimes.
@@ II Geodesic Completeness
[…]
@@
[…]
for null initial conditions. To solve the ODE we invert $\lambda\left(t\right)$, instead of using $t\left(\lambda\right)$; which is auspicious, because if Eqs. 2.4 and 2.5 diverge then the domain of $\gamma$ is cofinite and the geodesic ray has domain $\left[0,\infty\right)$ and is complete.
Eq. 2.3 defines a pregeodesic and thus has a geodesic reparameterization. Appropriately reparameterizing, there is a geodesic in $\Sigma$ which has $\beta$ as its image: thus, if $\left(\Sigma,g_{\Sigma}\right)$ is complete as a Riemannian manifold, then the limiting factor to geodesic completeness is Eq. 2.2 with solutions of Eqs. 2.2 and 2.3. The assumption of “generality” in a GFRW is the least restrictive logical assumption needed to have Eqs. 2.2 and 2.3 characterize geodesic completeness of GFRWs. Hence, the criteria for geodesic completeness of a GFRW spacetime is given by Thm. 2 of Lesnefsky et al. (2023b) (which builds on previous work Sanchez (1998)). The theorem states:
@@
[…]
@@ II.1 Singularities
Of course, it is possible for a spacetime to be geodesically incomplete even without encountering a curvature singularity. The most renowned example is the incompleteness encountered in the flat cosmological slicing of de Sitter space, where an observer following a geodesic into the past encounters a boundary in finite proper time (for a recent discussion see Kinney et al. (2023)).
@@
where $w$ is the equation of state parameter relating the pressure $p$ to the energy density $\rho$ of the fluid. We now force this fluid to violate the NEC for all time, thus also violating the ANEC. From the NEC condition $\rho+p=\rho(1+w)\geq 0$ (see Append. A), the fluid will have EOS parameter $w<-1$ in order to violate NEC. We take a flat $k=0$ FRW metric, and since $\dot{H}\propto-\rho(1+w)$, we find NEC violation implies $\dot{H}>0$ for $\rho>0$. We now take $\dot{H}=\ddot{a}/a-(\dot{a}/a)^{2}=2/\beta^{2}$, for a positive constant $\beta$, and integrate to get
@@
which diverges in a singular fashion as $t\rightarrow\infty$. The Kretschmann scalar $K=R_{\mu\nu\kappa\sigma}R^{\mu\nu\kappa\sigma}\propto t^{4}$, likewise diverges at $t\rightarrow\infty$. Geometrically, the geodesically connected domains [^6] For an FRW spacetime, the observable universe (for a given observer) is the geodesically connected domain of said observer at that point. The geodesically connected domain of a point $p\in\mathcal{M}$ consists of all points which can be reached by a geodesic. In contrast to a geodesically complete Riemannian manifold, where any two points are guaranteed to be connected by a geodesic, many geodesically complete Lorentzian manifolds are not geodesically connected. of any point become arbitrarily small as the scale factor and curvatures become arbitrarily large. Eventually, the geodesically connected domain becomes smaller than a Planck volume, although the affine parameter of a geodesic is still defined. Hence, in general, it is difficult to construct a physically reasonable eternal cosmological model with such extended period of NEC violation.
@@ II.2 Inflationary and bouncing cosmologies
Let $\left(\mathcal{M},g\right)$ be an $n$ dimensional spacetime which admits a connected open neighborhood $U\subset\mathcal{M}$ which is isometric to $\left(b,c\right)\times_{a}V$ as a warped product open submanifold, where $\left(b,c\right)$ is a timelike codimension $n-1$ embedded submanifold, $V$ is a connected spacelike codimension 1 embedded submanifold and no assumptions are made concerning $a(t)$ except that it is a well defined function between sets. The spacetime $\left(\mathcal{M},g\right)$ is an inflationary spacetime if there exists some $t_{0}\in\left(b,c\right)$ such that (assuming that it exists and is well defined), $\ddot{a}\left(t_{0}\right)>0$ where derivatives of the scale factor are taken with respect to the timelike $\left(b,c\right)$ coordinate and the sign of $\ddot{a}$ is given with respect to the time orientation of $\mathcal{M}$.
Note, in the above definition of inflation we make no assumptions about the continuity of the scale factor $a$ or its derivatives. Furthermore, this definition includes arbitrarily short periods of accelerated expansion as inflationary. We are aware some readers will find this objectionable; however, if one wishes to distinguish the terms “short accelerated expansion” from “inflation”, one is obliged to set an arbitrary scale for the duration of the acceleration. While GUT-scale inflation naturally requires at least 60 e-foldings of accelerated expansion to solve the flatness and horizon problems, low-energy (e.g., TeV-scale) inflation does the job with far fewer. It is only a few e-foldings of accelerated expansion that are actually probed in the cosmic microwave background radiation (CMB). In the present context of an eternal universe the difference between a few e-folding, and 60 e-foldings is entirely insignificant. For these reasons we find it unnecessary to reserve the term inflation to designate significant prolonged expansion and use the term interchangeably with accelerated expansion as is appropriate considering Defn. 1. This is the definition of “inflation” from Lesnefsky et al. (2023b), which we have reproduced here to align with our present notation.
@@
Let $\left(\mathcal{M},g\right)$ be an $n$ dimensional spacetime which admits a connected open neighborhood $U\subset\mathcal{M}$ which is isometric to $\left(b,c\right)\times_{a}V$ as a warped product open submanifold, where $\left(b,c\right)$ is a timelike codimension $n-1$ embedded submanifold, $V$ is a connected spacelike codimension 1 embedded submanifold and no assumptions are made concerning $a(t)$ except that it is a well defined non-constant function between sets. The spacetime $\left(\mathcal{M},g\right)$ is a bouncing spacetime if there exists some $t_{0}\in\left(b,c\right)$ such that (assuming that it exists and is well defined), $\dot{a}(t_{0})=0$, where derivatives of the scale factor are taken with respect to the timelike $\left(b,c\right)$ coordinate and the sign of $\dot{a}$ is given with respect to the time orientation of $\mathcal{M}$.
As we shall discover, in some cases it is advantageous to discuss a “bounce at infinity”. Considering that the point at infinity is strictly not an element [^8] One solution to this is to compactify. There are many compactification prescriptions, but typically one utilizes the Hausdorff single point compactification of $\mathbb{R}\hookrightarrow S^{1}$. Note, however, that Hausdorff single point compactification does not possess the universal property of preserving continuity upon function extension. of $\mathbb{R}$, we suggest:
@@
A neighborhood $\left(b,c\right)\times_{a}V\subset\left(\mathcal{M},g\right)$ is an open submanifold which is isometric to a GFRW $\left(b,c\right)\times_{a}\Sigma\subset\mathcal{M}$ is said to admit a bounce at positive infinity +$\infty$+ ∞ if $c\rightarrow+\infty$ is unbounded above and there exists a divergent cofinite sequence $\left\{t_{\ell}\right\}_{\ell=1}^{\infty}\rightarrow+\infty$ with $t_{k}>t_{\ell}$ for $k>\ell$ where $\left\{\dot{a}\left(t_{\ell}\right)\right\}_{\ell=1}^{\infty}\rightarrow 0$ constitutes a convergent Cauchy sequence.
A neighborhood $\left(b,c\right)\times_{a}V\subset\left(\mathcal{M},g\right)\subset\mathcal{M}$ is an open submanifold which is isometric to a GFRW which admits a bounce at negative infinity $\infty$- ∞ if $b\rightarrow-\infty$ is unbounded below and there exists a divergent cofinite sequence $\left\{t_{\ell}\right\}_{\ell=1}^{\infty}\rightarrow-\infty$ with $t_{k}<t_{\ell}$ for $k>\ell$ where $\left\{\dot{a}\left(t_{\ell}\right)\right\}_{\ell=1}^{\infty}\rightarrow 0$ constitutes a convergent Cauchy sequence.

### v2

@@ Abstract
We consider the possibility of a past and future eternal universe, constructing geodesically complete inflating, loitering, and bouncing spacetimes. We identify the constraints energy conditions in General Relativity place on the building of eternal cosmological models. Inflationary and bouncing behavior are shown to be essential ingredients in all significant examples. Non-trivial complete spacetimes are shown to violate the null energy condition (NEC) for at least some amount of time, as well as the average null energy condition unless curvature is included. Ignoring the intractable subtleties introduced by quantum considerations, such as rare tunneling events and Boltzmann brains, we demonstrate that these universes need not have had a beginning or an end.
@@ I Introduction
Whether or not the universe had a beginning is a question that naturally arises within our most successful theory of the early universe, the theory of inflation Guth (1981); Linde (1982); Albrecht and Steinhardt (1982). Soon after the theory’s inception it became clear that once inflation started, it would never stop. While early-time inflation has ceased in our observable universe, allowing galaxies and structures to form, it continues indefinitely in other parts of the universe, and hence, inflation is eternal into the future. Specifically, quantum fluctuations in the field responsible for the accelerated expansion can cause the field to remain in its high-energy state, sustaining inflation in those regions indefinitely Vilenkin (1983); Linde (1986); Guth (2007). In a Letter by Borde, Guth and Vilenkin (BGV) Borde et al. (2003), it was subsequently argued that inflation cannot likewise be eternal into the past; although, we have countered this belief Easson and Lesnefsky (2024); Lesnefsky et al. (2023) and present further evidence that inflation can, in fact, be eternal into the past in the current work. Acknowledging the possibility of a past eternal universe opens entirely new avenues of research. We show there are multiple ways that the universe can sustain itself within classical General Relativity for an eternity, albeit at the expense of violation of certain energy conditions.
Three broad categories of an eternal universe are suggested. Our examples are admittedly toy models; however, given the wide variety we expect our findings to be quite general and applicable to many solutions of interest. In particular, we will construct eternal inflationary, loitering, and bouncing cosmologies. In each case we will provide a rigorous proof of full geodesic completeness. Existing proposed no-go theorems are circumvented or obviated as discussed in Easson and Lesnefsky (2024). We show that all (non-trivial) models that are eternal geodesically complete necessarily involve a phase of accelerated expansion (or inflationary phase) for at least some amount of time and generically include a bouncing phase in accordance with our reported definitions.
The models we present violate the traditional energy conditions, a property we conjecture is necessary in GR for geodesic completeness.  [^3] A contained discussion of energy conditions is offered in Appendix A. In some instances violations of the null energy condition (NEC) can be limited to an arbitrarily short amount of time $\Delta t$, and confined to the primordial past;  [^4] In its most provocative form, the Heisenberg uncertainty principle is written $\Delta E\Delta t\geq\hbar/2$. If the time interval is made very small the uncertainty principle may result in a very large energy fluctuation which can back-react on the spacetime. Although, the models presented here ultimately violate the average null energy condition (ANEC) in a flat universe.
@@
$\int_{\gamma}T_{\mu\nu}k^{\mu}k^{\nu}d\lambda\geq 0\,$, (1.1)
where the integral is over a null geodesic $\gamma$ with tangent vector $k^{\mu}=dx^{\mu}/d\lambda$, and $\lambda$ is an affine parameter with respect to which the tangent vector to the geodesic is defined. In GR, it is frequently convenient to replace $T_{\mu\nu}$ in Eq. 1.1 by the Ricci tensor $R_{\mu\nu}$, simply using the Einstein field equations and the definition of a null vector. The ANEC has been shown to hold even in the case of arbitrary Casimir systems as long as the geodesic does not intersect or asymptotically approach the plates Graham and Olum (2005); Fewster et al. (2007). If the null geodesic is not assumed to be achronal, the ANEC may be violated by a quantum scalar field in a spacetime compactified in one spatial dimension or in the spacetime around a Schwarzschild black hole Visser (1996).
@@
A few of the Theorems and Corollaries we discuss were presented without proof in our earlier work Lesnefsky et al. (2023); Easson and Lesnefsky (2024). The proofs are revealed presently.
An outline of this paper is as follows. In section II, we prove a theorem which definitively provides the criteria for geodesic completeness of generalized Friedmann-Robertson-Walker (GFRW) spacetimes. In section III, we discuss the definition of inflation and bouncing cosmologies and provide an example of an eternally inflating universe which is eternal and geodesically complete both into the past and future. In section IV, we provide examples of eternal bouncing cosmologies which are geodesically complete both into the past and future. In section V, we provide examples of loitering (and quasi-loitering) models which are eternal and geodesically complete both into the past and future. Finally, in section VI, we present a discussion and our conclusions. We provide appendices to discuss the energy conditions in General Relativity with respect to cosmology, to provide proof details, to define Generalized Friedmann-Robertson-Walker (GFRW) spacetimes, and to present a detailed discussion of geodesics for several models.
@@ II Geodesic Completeness
$\lambda\left(t\right)=\int^{t}\frac{a\left(\zeta\right)}{\sqrt{a^{2}\left(\zeta\right)+1}}d\zeta\,$, (2.4)
@@
$\lambda\left(t\right)=\int^{t}a\left(\zeta\right)d\zeta\,$, (2.5)
for null initial conditions. To solve the ODE we invert $\lambda\left(t\right)$, instead of using $t\left(\lambda\right)$; which is auspicious, because if Eqs. 2.4 and 2.5 diverge then the domain of $\gamma$ is infinite and the geodesic ray has domain $\left[0,\infty\right)$ and is complete.
Eq. 2.3 defines a pregeodesic and thus has a geodesic reparameterization. Appropriately reparameterizing, there is a geodesic in $\Sigma$ which has $\beta$ as its image: thus, if $\left(\Sigma,g_{\Sigma}\right)$ is complete as a Riemannian manifold, then the limiting factor to geodesic completeness is Eq. 2.2 with solutions of Eqs. 2.2 and 2.3. The assumption of “generality” in a GFRW is the least restrictive logical assumption needed to have Eqs. 2.2 and 2.3 characterize geodesic completeness of GFRWs. Hence, the criteria for geodesic completeness of a GFRW spacetime is given by Thm. 2 of Lesnefsky et al. (2023) (which builds on previous work Sanchez (1998)). The theorem states:
@@
    The spacetime $\mathcal{M}$ is future null complete iff $\int_{t_{0}}^{\infty}a\left(\zeta\right)d\zeta$ diverges for all $t_{0}\in\mathbb{R}$.
@@ II.1 Singularities
Of course, it is possible for a spacetime to be geodesically incomplete even without encountering a curvature singularity. The most renowned example is the incompleteness encountered in the flat cosmological slicing of de Sitter space, where an observer following a geodesic into the past encounters a boundary in finite proper time (for a recent discussion see Kinney et al. (2023); Geshnizjani et al. (2023)).
@@
where $w$ is the equation of state parameter relating the pressure $p$ to the energy density $\rho$ of the fluid. We now force this fluid to violate the NEC for all time, thus also violating the ANEC. From the NEC condition $\rho+p=\rho(1+w)\geq 0$ (see Append. A), the fluid will have EOS parameter w<1w<-1 in order to violate NEC. We take a flat $k=0$ FRW metric, and since $\dot{H}\propto-\rho(1+w)$, we find NEC violation implies $\dot{H}>0$ for $\rho>0$. We now take $\dot{H}=\ddot{a}/a-(\dot{a}/a)^{2}=2/\beta^{2}$, for a positive constant $\beta$, and integrate to get
@@
which diverges in a singular fashion as $t\rightarrow\infty$. The Kretschmann scalar $K=R_{\mu\nu\kappa\sigma}R^{\mu\nu\kappa\sigma}\propto t^{4}$, likewise diverges at $t\rightarrow\infty$. Geometrically, the geodesically connected domains [^6] For an FRW spacetime, the observable universe (for a given observer) is the geodesically connected domain of said observer at that point. The geodesically connected domain of a point $p\in\mathcal{M}$ consists of all points which can be reached by a geodesic. In contrast to a geodesically complete Riemannian manifold, where any two points are guaranteed to be connected by a geodesic, many geodesically complete Lorentzian manifolds are not geodesically connected. of any point become arbitrarily small as the scale factor and curvatures become arbitrarily large. Eventually, the geodesically connected domain becomes smaller than a Planck volume, although the affine parameter of a geodesic remains well defined. Hence, in general, it is difficult to construct a physically reasonable eternal cosmological model with such extended period of NEC violation.
@@ II.2 Inflationary and bouncing cosmologies
Let $\left(\mathcal{M},g\right)$ be an $n$ dimensional spacetime which admits a connected open neighborhood $U\subset\mathcal{M}$ which is isometric to $\left(b,c\right)\times_{a}V$ as a warped product open submanifold, where $\left(b,c\right)$ is a timelike codimension n1n-1 embedded submanifold, $V$ is a connected spacelike codimension 1 embedded submanifold and no assumptions are made concerning $a(t)$ except that it is a well defined function between sets. The spacetime $\left(\mathcal{M},g\right)$ is an inflationary spacetime if there exists some $t_{0}\in\left(b,c\right)$ such that (assuming that it exists and is well defined), $\ddot{a}\left(t_{0}\right)>0$ where derivatives of the scale factor are taken with respect to the timelike $\left(b,c\right)$ coordinate and the sign of $\ddot{a}$ is given with respect to the time orientation of $\mathcal{M}$.
Note, in the above definition of inflation we make no assumptions about the continuity of the scale factor $a$ or its derivatives. Furthermore, this definition includes arbitrarily short periods of accelerated expansion as inflationary. We are aware some readers will find this objectionable; however, if one wishes to distinguish the terms “short accelerated expansion” from “inflation”, one is obliged to set an arbitrary scale for the duration of the acceleration. While GUT-scale inflation naturally requires at least 60 e-foldings of accelerated expansion to solve the flatness and horizon problems, low-energy (e.g., TeV-scale) inflation does the job with far fewer. It is only a few e-foldings of accelerated expansion that are actually probed in the cosmic microwave background radiation (CMB). In the present context of an eternal universe the difference between a few e-folding, and 60 e-foldings is entirely insignificant. For these reasons we find it unnecessary to reserve the term inflation to designate significant prolonged expansion and use the term interchangeably with accelerated expansion as is appropriate considering Defn. 1. This is the definition of “inflation” from Lesnefsky et al. (2023), which we have reproduced here to align with our present notation.
@@
Let $\left(\mathcal{M},g\right)$ be an $n$ dimensional spacetime which admits a connected open neighborhood $U\subset\mathcal{M}$ which is isometric to $\left(b,c\right)\times_{a}V$ as a warped product open submanifold, where $\left(b,c\right)$ is a timelike codimension n1n-1 embedded submanifold, $V$ is a connected spacelike codimension 1 embedded submanifold and no assumptions are made concerning $a(t)$ except that it is a well defined non-constant function between sets. The spacetime $\left(\mathcal{M},g\right)$ is a bouncing spacetime if there exists some $t_{0}\in\left(b,c\right)$ such that (assuming that it exists and is well defined), $\dot{a}(t_{0})=0$, where derivatives of the scale factor are taken with respect to the timelike $\left(b,c\right)$ coordinate and the sign of $\dot{a}$ is given with respect to the time orientation of $\mathcal{M}$.
We would like to highlight the fact that, in the the usual consideration of a universe shrinking to a turning point and then enlarging, Defn. 2 also considers an expanding universe reaching a turning point and then contracting a bounce as well.
As we shall discover, in some cases it is advantageous to discuss a “bounce at infinity”. Considering that the point at infinity is strictly not an element [^8] One solution to this is to compactify. There are many compactification prescriptions, but typically one utilizes the Hausdorff single point compactification of $\mathbb{R}\hookrightarrow S^{1}$. Note, however, that a Hausdorff single point compactification does not possess the universal property of preserving continuity upon function extension. of $\mathbb{R}$, we suggest:
@@
A neighborhood $\left(b,c\right)\times_{a}V\subset\left(\mathcal{M},g\right)$ is an open submanifold which is isometric to a GFRW $\left(b,c\right)\times_{a}\Sigma\subset\mathcal{M}$ is said to admit a bounce at positive infinity $\infty$ if, upon extension if necessary, $c\rightarrow+\infty$ is unbounded above and there exists a divergent sequence $\left\{t_{\ell}\right\}_{\ell=1}^{\infty}\rightarrow+\infty$ with $t_{k}>t_{\ell}$ for $k>\ell$ where $\left\{\dot{a}\left(t_{\ell}\right)\right\}_{\ell=1}^{\infty}\rightarrow 0$ constitutes a convergent Cauchy sequence for non-constant $a\left(t\right)$.
A neighborhood $\left(b,c\right)\times_{a}V\subset\left(\mathcal{M},g\right)\subset\mathcal{M}$ is an open submanifold which is isometric to a GFRW which admits a bounce at negative infinity -$\infty$ if, upon extension if necessary, $b\rightarrow$-$\infty$ is unbounded below and there exists a divergent sequence $\left\{t_{\ell}\right\}_{\ell=1}^{\infty}\rightarrow$-$\infty$ with $t_{k}<t_{\ell}$ for $k>\ell$ where $\left\{\dot{a}\left(t_{\ell}\right)\right\}_{\ell=1}^{\infty}\rightarrow 0$ constitutes a convergent Cauchy sequence.